a) Show that \sqrt{2}+\sqrt{3}=\sqrt{5+\sqrt{24}}

b) For non zero integers a,b,c,d exists in \mathbb{Z} investigate the conditions on a,b,c,d that are equivalent to the condition \sqrt{a}+\sqrt{b}= \sqrt{c+\sqrt{d}}
Answer #1

J67520 = 2+3+34x2xp = J 29+(33)*+852.08) - Sztole = 52+13 £ 55+520 = 52+ ④ Bat & Jetta)? => ) +(B) Teva 5$ = tud => a+olte Ja